# Generic consistency and nondegeneracy of vertically parametrized systems

### v3

@@ 3.7. Extension to $\mathbb{C}$
    Suppose that $(\mathcal{A}\times\mathcal{B},\mathcal{X})\subseteq\mathbb{C}^{m}\times\mathbb{C}^{\ell}\times\mathbb{C}^{n}$ is an algebraically defined pair, and assume that $\mathcal{I}_{\mathbb{C}}\cap(\mathcal{A}\times\mathcal{B}\times\mathcal{X})$ is Zariski dense in $\mathcal{I}_{\mathbb{C}}$. Then the affine analogs of (deg1), (degX1), (degXG), (degAllG), (setE), (setZ), (dimG), (dim1) are equivalent, any of these statements imply (dimX), (real) and (reg), and (rad) holds independently of the other statements.
@@
    For generic $(a,b)\in\mathbb{C}^{m}\times\mathbb{C}^{\ell}$, the variety $\mathbb{V}_{\mathbb{C}}(F_{a,b})$ has no irreducible components contained in the coordinate hyperplanes of $\mathbb{C}^{n}$.
@@
In the freely parametrized setting, we recover [LSEDSV21, Prop. 2.1].
@@
    There exists a linearly independent tuple $(v_{1},\ldots,v_{s})\in\prod_{i=1}^{s}\mathcal{S}_{i}$.

### v4

@@ 3.7. Extension to $\mathbb{C}$
    Suppose that $(\mathcal{A}\times\mathcal{B},\mathcal{X})\subseteq\mathbb{C}^{m}\times\mathbb{C}^{\ell}\times\mathbb{C}^{n}$ is an algebraically defined pair, and assume that $\mathcal{I}_{\mathbb{C}}\cap(\mathcal{A}\times\mathcal{B}\times\mathcal{X})$ is Zariski dense in $\mathcal{I}_{\mathbb{C}}$. Then the affine analogs of (deg1), (degX1), (degXG), (degAllG), (setE), (setZ), (flatG), (dim1), (dimG) are equivalent, any of these statements imply (dimX), (real) and (reg), and (rad) holds independently of the other statements.
@@
    For generic $(a,b)\in\mathbb{C}^{m}\times\mathbb{C}^{\ell}$, the variety $\mathbb{V}_{\mathbb{C}}(F_{a,b})$ has no irreducible component contained in a coordinate hyperplane of $\mathbb{C}^{n}$.
@@
In the freely parametrized setting, we recover [LSEDSV21, Proposition 2.1].
@@
    The tuple $(\mathcal{S}_{1},\ldots,\mathcal{S}_{s})$ admits an independent transversal.